increase of $25,000 after five years. He can also invest his money in risky assets where there are three outcomes: he has a 31 percent chance of losing $30,000, a 12 percent chance of gaining S15,000 and a 57 percent chance of gaining an unknown amount of money. What must the unknown amount of money equal if the expected value is to equal the guaranteed $25,000? Give your answer to the nearest whole dollar (with no decimal points, spaces, S signs, or commas in your answer)

Ans: 58018 Explanation Let the unknown amount is X Expected value -0.31(30,000) 0.12(15,000) 0.57X) 25,000 -9,300 1,800+0.57X 0.57X 25,0009,300 1,800 32,500 X 32,500/0.57- 57018 Thus, the unknown amount is 57018

Is this answer correct? What I calculated is not the same as this answer.
Answer #1

Answer

Expected Value of the gain = P1V1 + P2V2 + P3V3

Here, P1 = 0.31 + P2 = 0.12 and P3 = 0.57

V1 = Gain with probability P1 = -30,000 (-ve sign indicates that there is a loss instead of gain)

V2 = Gain with probability P2 = 15,000

V3 = Gain with probability P3 = X

Hence, P1V1 + P2V2 + P3V3 = 0.31*(-30000) + 0.12*15000 + 0.57*X

We want this equal to 25000

=> 0.31*(-30000) + 0.12*15000 + 0.57*X = 25000

=> 0.57X = 25000 - (0.31*(-30000) + 0.12*15000)

=> X = 57018

Hence, what you have done is absolutely correct.